OpenStudy (anonymous):

Two solenoids stand side-by-side. Solenoid B differs from solenoid A in three respects. Its total current is 2 times as great, it has 4 times as many coils, and it is 3 times as long. How do the magnetic fields inside the two solenoids compare? A. A standing wave is generated. B. Beats are produced. C. A traveling wave is generated. D. Resonance occurs. E. The frequency of the resultant wave decreases.

OpenStudy (anonymous):

are u typing?

OpenStudy (anonymous):

what are the options you are giving? Are u sure they are connected with this question? Bcoz' you are asking just to find the ratio of magnetic fields. Magnetic field inside a solenoid is given by B=uNi/l where u is permeability of the core of the solenoid. In air,it is 4π *10^-7 N=total no of turns of solenoid i=current l=length of solenoid You can compare all these values for both solenoids and hence u can find the ratio of B in both solenoids.

OpenStudy (anonymous):

A standing wave is generated. B. Beats are produced. C. A traveling wave is generated. D. Resonance occurs. E. The frequency of the resultant wave decreases.

OpenStudy (anonymous):

i fail to see any connection between the question u are asking and the options u are giving. However u are also asking the comparison of magnetic fields inside both the solenoids which i have shown how to do in my previous reply.

OpenStudy (anonymous):

oh my bad lol I just realized that

OpenStudy (anonymous):

A. Solenoid B’s magnetic field is 24 times greater than solenoid A's magnetic field. B. Solenoid A’s magnetic field is 24 times greater than solenoid B's magnetic field. C. Solenoid B’s magnetic field is 6 times greater than solenoid A's magnetic field. D. Solenoid A’s magnetic field is 2.7 times greater than solenoid B's magnetic field. E. Solenoid B’s magnetic field is 2.7 times greater than solenoid A's magnetic field.

OpenStudy (anonymous):

that's better. Can u find the answer now?

OpenStudy (anonymous):

im kinda confused on how to type it into the equation

OpenStudy (anonymous):

magnetic field in A =Ba =u*Na*ia/la where Na=no of turns in solenoid A ia=current in A la=length of A similarly, magnetic field in B =Bb =u*Nb*ib/lb where Nb=no of turns in solenoid B ib=current in B lb=length of B Hence Ba/Bb =(Na/Nb) *(ia/ib)*(lb/la) Now u are given Na/Nb=1/4 ia/ib =1/2 la/lb =1/3 Putting these value we get Ba/Bb =3/8 which means Bb is 2.7 times greater than Ba (option D)

OpenStudy (anonymous):

thankyou so much you are awesome

OpenStudy (anonymous):

im in this class right now cause I failed physics lol

OpenStudy (anonymous):

that's fine. Feel free to ask!